OpenStudy (anonymous):

A rectangular playground is enclosed on three sides by a fence and one side of a house. The length of the fence is 20 yards. Given that the width of the playground is x yards, find: an expression for the lenth og the playground in temrs of x the area of the playground in terms of x the actual length and width of the playground that will give the maximum area

OpenStudy (anonymous):

length of playground + 2x = 20 so length of playground = 20 - 2x area = A = (20 - 2x)x = 20x - 2x^2 dA/dx = 20 - 4x dA/dx = 0 when x = 5 thus width = 5, length = 10 maximizes the area (maximum area = 50)

OpenStudy (anonymous):

or you could find maximum area using the turning point: A = 20x - 2x^2 = -2x^2 + 20x = -2(x^2 - 10x) = -2((x - 5)^2 - 25) = -2(x - 5)^2 + 50